# Generalizing Graph Transformers Across Diverse Graphs and Tasks via Pre-Training on Industrial-Scale Data

### v3

@@ Abstract.
Graph pre-training has been concentrated on graph-level on small graphs (e.g., molecular graphs) or learning node representations on a fixed graph. Extending graph pre-trained models to web-scale graphs with billions of nodes in industrial scenarios, while avoiding negative transfer across graphs or tasks, remains a challenge. We aim to develop a general graph pre-trained model with inductive ability that can make predictions for unseen new nodes and even new graphs. In this work, we introduce a scalable transformer-based graph pre-training framework called PGT (Pre-trained Graph Transformer). Specifically, we design a flexible and scalable graph transformer as the backbone network. Meanwhile, based on the masked autoencoder architecture, we design two pre-training tasks: one for reconstructing node features and the other one for reconstructing local structures. Unlike the original autoencoder architecture where the pre-trained decoder is discarded, we propose a novel strategy that utilizes the decoder for feature augmentation. We have deployed our framework on Tencent’s online game data. Extensive experiments have demonstrated that our framework can perform pre-training on real-world web-scale graphs with over 540 million nodes and 12 billion edges and generalizes effectively to unseen new graphs with different downstream tasks. We further conduct experiments on the publicly available ogbn-papers100M dataset, which consists of 111 million nodes and 1.6 billion edges. Our framework achieves state-of-the-art performance on both industrial datasets and public datasets, while also enjoying scalability and efficiency.
Graph Transformers, Self-Supervised Learning, Graph Pre-Training
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Neural networksccs: Information systems Social networks
1. Introduction
Online gaming, as a domain of interest in network science, presents a rich source of data due to its inherent graph-like structure. With a massive user base of 800 million online gaming users [^1] https://www.tencent.com/en-us/articles/2200928.html, Tencent is an influential player in the global gaming landscape.
@@ 1. Introduction
Challenges. While there have been some initial attempts to apply graph neural networks in real-world industrial settings (Wang et al., 2019; Zhang et al., 2023), deploying GNNs in products and bridging the gap between academia and industry remains a long journey.
@@
Firstly, message-passing GNNs, constrained by strong inductive biases, do not consistently perform well on real-world industrial graph data. Graph data in the industry may unavoidably contain structural noise, such that GNNs, which aggregate information from pre-defined neighbors, learn indistinct representations (Fox and Rajamanickam, 2019).
@@
Secondly, existing graph learning methods often struggle to meet the requirements of scalability and efficiency in inference posed by online gaming businesses. In addition, existing scalable GNNs tend to prioritize training over inference (Chiang et al., 2019; Zeng et al., 2019; Zhang et al., 2021a, b; He and Ma, 2022). This leads to a need to design novel sampling methods to meet the business demands for high-throughput and low-latency model inference.
@@
Thirdly, differing from most existing works that focus on optimizing models for a specific task on a fixed graph, the online gaming industry often involves highly diverse graphs and downstream tasks. Training models from scratch for each task on each graph can result in a significant waste of time and computational resources. Given the great success of pre-trained models in computer vision (Radford et al., 2021) and natural language processing (Brown et al., 2020; Sui et al., 2024), intuitively, we aim for a graph pre-trained model that serves as a general model that can effectively generalize across different graphs and tasks. However, in practice, due to the high diversity of graph data, graph pre-trained models often exhibit negative transfer, which hinders their performance (Hu et al., 2019; Qiu et al., 2020; Cao et al., 2023).
Contributions. In this work, we present PGT, a new graph pre-training framework with transformers. We make critical designs in various aspects, including backbone networks, sampling algorithms, and pre-training tasks.
@@
Instead of message-passing GNNs, we adopt the Transformer model (Vaswani et al., 2017) as the backbone network, which has demonstrated its superiority in various domains (Devlin et al., 2018; Dosovitskiy et al., 2021). Graph transformers disregard pre-defined graph structures and learn a soft, fully connected graph structure bias across all nodes. It can mitigate the graph structural noise and heterophily present in industrial data.
Applying the Transformer architecture to web-scale graphs is not a trivial endeavor, we employ the classic graph algorithm Personalized PageRank (PPR) to sample contextual node sequences for each seed node. By sampling sequences that are short but informative, the seed nodes are enabled to aggregate information from neighbors that are most relevant to them in the graph structure. Experiments show our method accelerates inference by up to 12.9x while achieving better performance compared to previous methods.
@@
The objective of graph pre-trained models is to generalize to unseen new nodes, edges, and even entirely novel graphs. Based on the Masked Graph Modeling (MGM) paradigm, we design two distinct pre-training tasks: one at the node-level for node feature reconstruction, and the other at the edge-level for local structural reconstruction. In the inference phase, we introduce a novel feature augmentation technique. In contrast to the conventional Masked Autoencoder (MAE) architecture, which discards the decoder after pre-training (He et al., 2022), we utilize the pre-trained decoder to generate reconstructed node features.
To summarize, our work makes the following four contributions:
@@
  • We conduct extensive experiments on both industrial and public datasets, demonstrating that PGT can effectively generalize to diverse graphs and tasks. Our empirical findings substantiate that PGT yields superior results compared to other graph self-supervised learning algorithms.
2. Related work
Pre-training on graphs. The proliferation of GNNs has led to significant efforts being dedicated to pre-train GNNs. Hu et al. (Hu et al., 2019) first propose to pre-train GNNs on molecular graphs with supervised strategies. GROVER (Rong et al., 2020) and MoCL (Sun et al., 2021) design more supervised pre-training tasks and achieve success in biomedical domains.
Similarly to the NLP and CV fields, more efforts focus on designing self-supervised pre-training strategies. Contrastive learning on graphs aims to learn representations by maximizing the mutual information between positive samples while minimizing it between negative samples. DGI (Veličković et al., 2018) and MVGRL (Hassani and Khasahmadi, 2020) design cross-scale contrast between node-level and graph-level representations. GRACE (Zhu et al., 2020), GCA (Zhu et al., 2021), and GraphCL (Hafidi et al., 2022) adopt different data augmentation methods to generate same-scale contrastive views. Following BYOL (Grill et al., 2020), BGRL (Thakoor et al., 2021) uses an additional target encoder to introduce self-distillation as a means to avoid negative sampling. Generative self-supervised learning on graphs, exemplified by GAE (Kipf and Welling, 2016b), employs an encoder-decoder architecture with the training objective of reconstructing input data to learn low-dimensional embeddings. Following in the footsteps of GAE, most generative learning methods take the original or perturbed graph structure and node features as input, designing tasks for feature reconstruction (e.g., MGAE (Wang et al., 2017), GraphMAE (Hou et al., 2022), GraphMAE2 (Hou et al., 2023), GMAE (Zhang et al., 2022a), and UniGraph (He and Hooi, 2024)) or structure reconstruction (e.g., VGAE (Kipf and Welling, 2016b), MaskGAE (Li et al., 2023)).
3. Preliminaries
@@
4. The PGT Framework
@@
Figure 1. Overview of PGT framework during pre-training phase. (1) Given a seed node $v_{s}$, we adopt the personalized PageRank (PPR) algorithm to sample a contextual node sequence $S_{s}$ to represent its local graph structure. (2) Subsequently, we randomly mask a subset of nodes on each sequence and feed the sequence composed of the unmasked nodes into the graph transformer encoder. The output consists of the embeddings of each unmasked node. (3) The decoding comprises two learning objectives: a) we incorporate the masked nodes into the output of the encoder and initialize them with a learnable token $[M]$. A shallow graph transformer is employed as a decoder to reconstruct the input features of the masked nodes. b) In the context of graphs, each sequence $S_{s}$ can be interpreted as a local neighborhood or subgraph. We adopt a simple MLP and contrastive loss aiming to make nodes within each sequence similar in the latent space while making them dissimilar to nodes from other sequences. This encourages the learned embeddings to reflect the actual connectivity patterns within local graph structures.
@@
In this section, we present our proposed PGT framework, which utilizes the MGM pretext task to pre-train graph transformers. The overall framework is illustrated in Figure 1.
4.1. Graph to Node Sequences
@@
Due to the quadratic computational and memory complexity of self-attention with respect to the input sequence length, we are unable to scale it to large-scale graphs, such as social networks. In this paper, we propose to sample contextual node sequences for each given seed node. Specifically, we utilize the personalized PageRank (PPR) algorithm to sample nodes containing the most relevant information for each seed node as the sequence input for the Transformer model. We define the following mathematical procedure: Given a graph represented as $G=(V,E)$, we consider a set of seed nodes $\bm{S}$ as input. The transition matrix $\bm{P}$ is constructed as $\bm{P}_{ij}=\frac{\bm{A}_{ij}}{\bm{d}_{i}}$, where $\bm{A}$ is the adjacency matrix of $G$, $\bm{d}_{i}$ is the out-degree of node $v_{i}$, and $\bm{P}_{ij}$ represents the probability of transitioning from node $v_{i}$ to node $v_{j}$. Next, we calculate the personalized PageRank scores for each node in the graph by solving the equation $\bm{r}=(1-\alpha)\bm{r}_{0}+\alpha\bm{P}^{T}\bm{r}$, where $\bm{r}$ is the personalized PageRank vector, $\bm{r}_{0}$ is the initial personalized vector with $\bm{r}_{0}(i)=1$ for $i$ in $\bm{S}$ and $\bm{r}_{0}(i)=0$ otherwise, and $\alpha$ is a damping factor. After obtaining the personalized PageRank scores, we select contextual nodes for each seed node $v_{s}$ based on the top-k nodes with the highest personalized PageRank scores:
$C_{s}=\{v_{s,i}\,|\,v_{s,i}\in\text{Top-}k\text{ nodes with highest }PPR(v_{s,i},v_{s})\}.$
@@
This sampling strategy offers several advantages: (1) It can be regarded as utilizing graph structural information to perform masking on full-graph attention, implicitly introducing inductive bias. (2) It unifies the sampling strategies for training and inference. Differing from sampling algorithms like GraphSAINT (Zeng et al., 2019) and GraphSAGE (Hamilton et al., 2017), which mainly focus on training and employ full neighborhood aggregation during inference, our method enables the use of the same node sequences during both the training and the inference phases. (3) It decouples the sampling of labeled nodes and auxiliary nodes. Whereas, our method only needs to extract the corresponding sequences, and the time cost scales with the number of labeled nodes.
4.2. Masked Encoding

### v4

@@ Abstract
Graph pre-training has been concentrated on graph-level tasks involving small graphs (e.g., molecular graphs) or learning node representations on a fixed graph. Extending graph pre-trained models to web-scale graphs with billions of nodes in industrial scenarios, while avoiding negative transfer across graphs or tasks, remains a challenge. We aim to develop a general graph pre-trained model with inductive ability that can make predictions for unseen new nodes and even new graphs. In this work, we introduce a scalable transformer-based graph pre-training framework called PGT (Pre-trained Graph Transformer) [^1] The code is available at https://github.com/yf-he/PGT.. Based on the masked autoencoder architecture, we design two pre-training tasks: one for reconstructing node features and the other for reconstructing local structures. Unlike the original autoencoder architecture where the pre-trained decoder is discarded, we propose a novel strategy that utilizes the decoder for feature augmentation. Our framework, tested on the publicly available ogbn-papers100M dataset with 111 million nodes and 1.6 billion edges, achieves state-of-the-art performance, showcasing scalability and efficiency. We have deployed our framework on Tencent’s online game data, confirming its capability to pre-train on real-world graphs with over 540 million nodes and 12 billion edges and to generalize effectively across diverse static and dynamic downstream tasks.
I Introduction
Graph data, characterized by its intricate network of interconnected nodes and edges, pervades a multitude of domains in the real world. This pervasive presence arises from the inherent relational structures observed in various systems, ranging from social networks [liao2018attributed, sui2024fidelis] and transportation networks [guo2019attention] to biological networks [ying2021transformers] and financial systems [wang2021review]. Graph Neural Networks (GNNs) have been a dominant model architecture in modeling complex relationships within graph data [kipf2016semi, velivckovic2017graph, hamilton2017inductive], making them versatile tools for addressing real-world challenges and fostering innovation [ying2018graph, masters2022gps++, allamanis2022graph, DBLP:journals/tkde/HuHQFX24, DBLP:conf/aaai/0016HHW25, sui2024can, liu2025efficient, liu2025guardreasoner, he2025enabling, he2025evotest, chen2025can, chen2025robustness, wang2025safety, he2025evaluating, zhang2021scr, he2022sgkd, sui2025meta].
Figure 1: Illustrations of graph pre-training in the online gaming industry.
Online gaming, as a domain of interest in network science, presents a rich source of data due to its inherent graph-like structure. Advanced graph learning techniques are increasingly recognized as an important means for enhancing data analysis in the online gaming industry [yang2022large, cheng2025lps].
@@ I Introduction
Challenges. While there have been some initial attempts to apply graph neural networks in real-world industrial settings [wang2019deep, zhang2023inferturbo], deploying GNNs in products and bridging the gap between academia and industry remains a long journey.
@@
Firstly, message-passing GNNs, constrained by strong inductive biases, do not consistently perform well on real-world industrial graph data. Graph data in the industry may unavoidably contain structural noise, causing GNNs, which aggregate information from pre-defined neighbors, to learn indistinct representations [fox2019robust]. An example is heterophilic graphs that do not adhere to the homophily assumption, where node features and labels significantly differ from those of their neighbors [zheng2022graph]. Designing an appropriate backbone network for online gaming networks is a significant challenge.
@@
Secondly, existing graph learning methods often struggle to meet the requirements of scalability and efficiency in inference posed by online gaming businesses. The size of real-world online gaming networks in industry is typically 5-6 times larger than the largest publicly available datasets in academia. In addition, existing scalable GNNs tend to prioritize training over inference [chiang2019cluster, zeng2019graphsaint]. This leads to a need to design novel sampling methods to meet the business demands for high-throughput and low-latency model inference.
@@
Thirdly, as illustrated in Fig. 1, differing from most existing works that focus on optimizing models for a specific task on a fixed graph, the online gaming industry often involves highly diverse graphs and downstream tasks. For example, friend relationships in various games can be modeled as different graphs, and within a single game, there exist various prediction tasks like user classification and friend recommendations. Training models from scratch for each task on each graph can result in a significant waste of time and computational resources. Given the great success of pre-trained models in computer vision [radford2021learning] and natural language processing [brown2020language], intuitively, we aim for a graph pre-trained model that serves as a general model that can effectively generalize across different graphs and tasks. However, in practice, due to the high diversity of graph data, graph pre-trained models often exhibit negative transfer, which hinders their performance [hu2019strategies, qiu2020gcc, cao2023pre, he2025unigraph, he2025unigraph2].
@@
Contributions. In this work, we present PGT, a new graph pre-training framework with transformers. Instead of message-passing GNNs, we adopt the Transformer model [vaswani2017attention] as the backbone, which has demonstrated its superiority in various domains [devlin2018bert, dosovitskiy2020image]. Graph transformers disregard pre-defined graph structures and learn a soft, fully connected graph structure bias across all nodes. This approach can mitigate the graph structural noise and heterophily present in industrial data.
Applying the Transformer architecture to web-scale graphs is not a trivial endeavor, primarily due to the computational and memory costs induced by the quadratic complexity of the self-attention mechanism. We employ the classic Personalized PageRank (PPR) to sample contextual node sequences for each seed node. This also decouples the sampling of labeled and auxiliary nodes, aligning well with industrial scenarios where labeled nodes constitute an extremely small fraction of the entire graph. In contrast to expensive random data accesses on the entire graph [chiang2019cluster, zeng2019graphsaint], our method merely accesses the corresponding node sequences. Experiments show our method accelerates inference by up to 12.9x while achieving better performance compared to previous methods.
@@
In the online gaming industry, due to the growing complexity and scale of games themselves and the interactions between players, entirely new graphs will continue to emerge as the business grows. The objective of graph pre-trained models is to generalize to unseen new nodes, edges, and even entirely novel graphs. The scalable graph transformer model we propose effectively mitigates the issue of structural distribution discrepancies between graphs (such as degree distribution) by implicitly learning the local structure of the graph from contextual sequences. Based on the Masked Graph Modeling (MGM) paradigm, we design two distinct pre-training tasks: one at the node-level for node feature reconstruction, and the other at the edge-level for local structural reconstruction. In the inference phase, we introduce a novel feature augmentation technique. In contrast to the conventional Masked Autoencoder (MAE) architecture, which discards the decoder after pre-training [he2022masked], we utilize the pre-trained decoder to generate reconstructed node features. These reconstructed features are then averaged with the original features and used as input for the encoder during inference.
We conduct extensive experiments on Tencent’s online gaming data to demonstrate the performance, efficiency, and transferability of PGT. We perform pre-training on a Tencent gaming user friendship graph comprising 540 million nodes and 12 billion edges. Subsequently, the pre-trained model is applied to two downstream tasks across four different games. Additionally, we experiment with publicly available benchmarks. The results on Tencent’s online gaming dataset and public datasets suggest that, when compared to existing graph self-supervised learning algorithms, the PGT framework is capable of achieving state-of-the-art performance.
Also, graph pre-training extends beyond online gaming applications and addresses fundamental challenges in graph learning across various domains. In finance, graph pre-trained models can enhance fraud detection and credit risk assessment by capturing complex relationships between transactions and entities despite heterogeneous graph structures. In healthcare, these models can improve patient outcome prediction and drug discovery by transferring knowledge across different medical data graphs with varied structures. The scalability and cross-graph generalization capabilities of our approach are particularly valuable in these domains, where graph data often exhibits significant structural noise and heterophily. Through our comprehensive experiments on both gaming data and public academic datasets spanning different domains, we demonstrate the broad applicability of our framework across diverse real-world settings.
Overall, our work makes the following four contributions:
@@
  • We conduct extensive experiments on both industrial and public datasets, demonstrating that PGT can effectively generalize to diverse graphs and tasks. Crucially, we show that a simple adaptation of our statically pre-trained model can serve as a powerful backbone for dynamic graph tasks, outperforming specialized dynamic GNNs trained from scratch. Our empirical findings substantiate that PGT yields superior results compared to other graph self-supervised algorithms.
II Related work
II-A Scalable Graph Neural Networks
Standard Graph Neural Networks (GNNs) [hamilton2017inductive, kipf2016semi, velivckovic2017graph, xu2018powerful] struggle with large-scale graphs due to memory constraints. A primary solution is sampling, including node-level sampling (GraphSAGE [hamilton2017inductive]) and sub-graph sampling (Cluster-GCN [chiang2019cluster], GraphSAINT [zeng2019graphsaint]). While most work focuses on training, methods like ShaDow-GNN [zeng2021decoupling] also address scalable inference. A complementary approach simplifies the GNN architecture itself. Linear GNNs like SGC [wu2019simplifying] remove non-linearities, collapsing propagation into a single matrix operation. Other graph reduction methods use techniques like coarsening or sparsification to shrink the graph [ma2025acceleration]. However, these simplification techniques can lose crucial structural information on large, heterogeneous industrial graphs.
II-B Graph Transformers
Inspired by their success in NLP [vaswani2017attention] and computer vision [dosovitskiy2020image], Transformers have been adapted for graph learning. Early Graph Transformers focused on small graphs, incorporating structural information via positional encodings, such as Laplacian eigenvectors (GT [dwivedi2020generalization], LiteGT [chen2021litegt], SAN [chen2022structure]) or centrality measures (Graphormer [ying2021transformers]). The quadratic complexity of self-attention hinders their application to large graphs. To address this, scalable Graph Transformers employ techniques like neighbor sampling (Gophormer [zhao2021gophormer], ANS-GT [zhang2022hierarchical]) or efficient attention approximations like kernelized operators (NodeFormer [wu2022nodeformer]).
II-C Pre-training on Graphs.
Pre-training has become a key paradigm for GNNs. Initial work used supervised pre-training on domain-specific data [hu2019strategies]. More recent efforts focus on self-supervised learning, which broadly falls into two categories. Contrastive methods learn by maximizing agreement between positive views, using cross-scale contrast (DGI [velivckovic2018deep], MVGRL [hassani2020contrastive]) or, inspired by BYOL [grill2020bootstrap], avoiding negative samples via self-distillation (BGRL [thakoor2021bootstrapped]). Generative methods use an encoder-decoder framework to reconstruct masked node features (MGAE [wang2017mgae], GraphMAE [hou2022graphmae], GraphMAE2 [hou2023graphmae2], GMAE [zhang2022graph]) or graph structure (GAE [kipf2016variational], VGAE [kipf2016variational], MaskGAE [li2022maskgae]). However, the scalability and cross-graph transferability of these methods for web-scale industrial graphs with varying structures remain open challenges [qiu2020gcc, cao2023pre].
II-D Learning on Dynamic Graphs
A related but distinct challenge is learning on dynamic graphs, where the graph structure and features evolve over time. While our work focuses on pre-training for generalization across diverse static graphs, dynamic graph representation learning methods are specifically designed to capture temporal patterns. These models, such as EvolveGCN [pareja2020evolvegcn] and TGN [rossi2020temporal], often use mechanisms like recurrent networks or memory modules to efficiently update representations as the graph changes. DGCN [gao2022novel] utilizes an LSTM to update the weight parameters of a GCN, aiming to capture global structural information across temporal snapshots. Other notable approaches include VGRNN [hajiramezanali2019variational], a variational generative model. Although some of our industrial tasks involve temporal snapshots, our core contribution addresses the orthogonal problem of negative transfer between structurally different graphs, rather than modeling temporal evolution within a single graph.
III Preliminaries
@@
III-A Graph Pre-training: Challenges
The goal of graph pre-training is to learn a model on a graph domain that generalizes to unseen graphs and benefits various downstream tasks. Achieving this presents two primary challenges. First, the model’s backbone must have sufficient capacity to learn transferable knowledge while remaining scalable for both training and inference. Many existing scalable GNNs are shallow and focus primarily on training efficiency, often overlooking inference [chiang2019cluster, zeng2019graphsaint]. Second, and more critically, pre-trained models often suffer from negative transfer when applied across graphs with diverse structures [hu2019strategies, qiu2020gcc]. For instance, social networks can vary from dense, bidirectional graphs (like Facebook) to sparser, star-shaped structures (like Twitter). The strong inductive biases of GNNs make it difficult to generalize across such divergent local structures.
Existing graph neural networks exhibit specific limitations when confronted with diverse graph structures, primarily due to their rigid message-passing mechanisms. Message-passing GNNs inherently assume that connected nodes share similar features or labels (homophily), which proves problematic in heterophilic graphs where connected nodes tend to have dissimilar characteristics. For instance, in online gaming networks, users with opposing play styles or from different demographic groups often connect, creating heterophilic patterns that conventional GNNs struggle to model. Additionally, these models are highly sensitive to structural noise—common in industrial graphs due to frequent changes in user relationships, spam connections, or adversarial behaviors. For example, in online gaming networks, temporary team formations or automated friend requests can introduce misleading structural patterns. This issue is particularly severe in cross-graph transfer scenarios, where pre-trained GNNs fail to generalize because they overly adapt to the specific structural patterns of the training graph. Empirical studies show that when deployed on graphs with different structural characteristics (e.g., degree distributions, clustering coefficients, or assortativity), pre-trained GNNs often perform worse than randomly initialized models, demonstrating negative transfer effects.
III-B Negative Transfer in Graph Pre-training
Definition 1 (Negative Transfer in Graph Pre-training)
Let $M_{pre}$ be a graph model pre-trained on source graph $G_{s}$, and $M_{rand}$ be the same model architecture but with random initialization. Negative transfer occurs when, after applying both models to a downstream task on target graph $G_{t}$ (either through linear probing or fine-tuning), the performance of $M_{pre}$ is worse than $M_{rand}$, i.e., $\text{Perf}(M_{pre},G_{t})<\text{Perf}(M_{rand},G_{t})$, where $\text{Perf}(\cdot,\cdot)$ is a task-specific metric.
IV The PGT Framework
@@
Figure 2: Overview of PGT framework during pre-training phase. (1) Given a seed node $v_{s}$, we adopt the personalized PageRank (PPR) algorithm to sample a contextual node sequence $S_{s}$ to represent its local graph structure. (2) Subsequently, we randomly mask a subset of nodes in each sequence and feed the sequence composed of the unmasked nodes into the graph transformer encoder. The output consists of the embeddings of each unmasked node. (3) The decoding comprises two learning objectives: a) we incorporate the masked nodes into the output of the encoder and initialize them with a learnable token $[M]$. A shallow graph transformer is employed as a decoder to reconstruct the input features of the masked nodes. b) In the context of graphs, each sequence $S_{s}$ can be interpreted as a local neighborhood or subgraph. We adopt a simple MLP and contrastive loss aiming to make nodes within each sequence similar in the latent space while making them dissimilar to nodes from other sequences. This encourages the learned embeddings to reflect the actual connectivity patterns within local graph structures.
@@
In this section, we present our proposed PGT framework, which utilizes the MGM pretext task to pre-train graph transformers. The overall framework is illustrated in Fig. 2.
IV-A Graph to Node Sequences
@@
Due to the quadratic computational and memory complexity of self-attention with respect to the input sequence length, we are unable to scale it to large-scale graphs, such as social networks. In this paper, we propose to sample contextual node sequences for each given seed node. Specifically, we utilize the personalized PageRank (PPR) algorithm to sample nodes containing the most relevant information for each seed node as the sequence input for the Transformer model. We define the following mathematical procedure: Given a graph represented as $G=(V,E)$, we consider a set of seed nodes $\boldsymbol{S}$ as input. The transition matrix $\boldsymbol{P}$ is constructed as $\boldsymbol{P}_{ij}=\frac{\boldsymbol{A}_{ij}}{\boldsymbol{d}_{i}}$, where $\boldsymbol{A}$ is the adjacency matrix of $G$, $\boldsymbol{d}_{i}$ is the out-degree of node $v_{i}$, and $\boldsymbol{P}_{ij}$ represents the probability of transitioning from node $v_{i}$ to node $v_{j}$. Next, we calculate the personalized PageRank scores for each node in the graph by solving the equation $\boldsymbol{r}=(1-\alpha)\boldsymbol{r}_{0}+\alpha\boldsymbol{P}^{T}\boldsymbol{r}$, where $\boldsymbol{r}$ is the personalized PageRank vector, $\boldsymbol{r}_{0}$ is the initial personalized vector with $\boldsymbol{r}_{0}(i)=1$ for $i$ in $\boldsymbol{S}$ and $\boldsymbol{r}_{0}(i)=0$ otherwise, and $\alpha$ is a damping factor. After obtaining the personalized PageRank scores, we select contextual nodes for each seed node $v_{s}$ based on the top-k nodes with the highest personalized PageRank scores:
$C_{s}=\{v_{s,i}\,|\,v_{s,i}\in\text{Top-}k\text{ nodes with highest }\mathrm{PPR}(v_{s,i},v_{s})\}.$
@@
Figure 3: Step-by-step illustration of the PPR sampling process: (1) Original graph with seed node, (2) Computation of Personalized PageRank scores, (3) Selection of top-$k$ nodes, and (4) Final sequence construction for transformer input.
Figure 4: Comparison of PPR sampling for different tasks. Top row: Node classification task with a single seed node. Bottom row: Link prediction task with source and target nodes, showing how contexts are sampled from both endpoints.
As illustrated in Fig. 3, PPR sampling follows a four-step process: (1) starting with a seed node, (2) computing importance scores that balance both proximity and centrality, (3) selecting the top-$k$ nodes with highest scores, and (4) constructing an ordered sequence for transformer input. Unlike traditional distance-based sampling approaches that may select unimportant neighbors or miss influential nodes slightly further away, PPR naturally identifies nodes that are both structurally important and meaningfully connected to the seed node. This creates balanced contextual sequences that effectively represent the local graph topology. As demonstrated in Fig. 4, this sampling approach is adaptable across different downstream tasks: for node classification (top row), PPR samples contexts that help determine class membership, while for link prediction (bottom row), we sample from both endpoints of the potential edge, with the overlap between contexts providing valuable signal for prediction.
@@
This sampling strategy offers several advantages: (1) It can be regarded as utilizing graph structural information to perform masking on full-graph attention, implicitly introducing inductive bias. (2) It unifies the sampling strategies for training and inference. Unlike sampling algorithms like GraphSAINT [zeng2019graphsaint] and GraphSAGE [hamilton2017inductive], which mainly focus on training and employ full neighborhood aggregation during inference, our method enables the use of the same node sequences during both the training and the inference phases. (3) It decouples the sampling of labeled nodes and auxiliary nodes, as our method only needs to extract the corresponding sequences, and the time cost scales with the number of labeled nodes.
For pre-training, we randomly select 10% of all nodes as seed nodes to ensure computational efficiency while maintaining representative coverage of the graph. During inference for downstream tasks, the seed nodes are simply the nodes requiring predictions, which effectively decouples the sampling of labeled nodes (typically scarce) from auxiliary nodes, enhancing efficiency in industrial applications.
IV-B Masked Encoding